# Benchmarking Predictive Coding Networks – Made Simple

### v1

@@ 6 Computational Resources and Implementations Details
The efficiency of PCX could be further increased by fully parallelizing all the computations occuring within a PCN. In fact, in its current state, JIT seems to be unable to parallelize the executions of the layers; a problem that can be addressed with the JAX primitive vmap, but only in the unpractical case where all the layers have the same dimension. To test how different hyperparameters of the model influence the training speed, we have taken a feedforward model, and trained it multiple times, each time increasing a specific hyperparameter by a multiplicative factor. The results, reported in Fig. 8, show that the two parameters that increase the training time are the number of layers $L$ (when the calculations are not explicitly vmapped), and the number of steps $T$. Ideally, only $T$ should affect the training time as inference is an inherently sequential process that cannot be parallelized, but this is not the case, as the time scales linearly with‘F the amount of layers. Details are reported in Appendix F.
@@ 7 Discussion
The main contribution of this work is the introduction and open-source release that can be used to perform deep learning tasks using PCNs. Its efficiency relies on JAX’s Just-In-Time compilation and carefully structured primitives built to take advantage of it. A second advantage of our library is its intuitive setup, tailored to users already familiar with other deep learning frameworks such as PyTorch. Together with the large number of tutorials we release will make it easy for new users to train networks using PC.
We have also performed a large-scale comparison study on image classification and image generation tasks, unifying under the same computational framework multiple optimization algorithms for PCNs present in the literature. In addition, we have tried multiple, popular, optimizers and training techniques, as well as an extremely large choice of hyperparameters. For CIFAR100 only, for example, we have conducted thousands of individual experiments, that have been used to obtain new state of the art results, as well as provide insights on what works and what does not, that will be useful in the future to researchers tackling deep learning problem with PCNs.
@@ Acknowledgments
This work was supported by Medical Research Council grant MC_UU_00003/1 awarded to RB.
@@
The code for PCX is available at https://github.com/liukidar/pcax. The experiments conducted in the main body and in the appendix can be found in the examples folder.
@@ Appendix A PCX – A Brief Introduction
In Section 3, we defined PCNs as models with parameters $\theta=\{\theta_{0},\ldots,\theta_{L}\}$ and state $h=\{h_{0},\ldots,h_{L}\}$. In PCX, we divide a model in components of two main categories: layers (i.e., the traditional deep-learning transformations such as ’Linear’ or ’Conv2D’) and vodes (i.e., vectorized nodes that store the array of neurons representing state $h_{l}$). A PCN is defined as follows: {minted}python import jax.nn as jnn import pcx.predictive_coding as pxc import pcx.nn as pxnn
@@
return u In the __call__ method, we forward the input $x$ through the network. Note that every time we call a vode, we are effectively storing in it the activation $u_{l}$ (so that we can later compute the energy $\epsilon_{l}^{2}$ associated to the vode) and returning its state $h_{l}$ (i.e., x = vode(u) corresponds to vode.set("u", u); x = vode.get("h")). During training, the label $y$ is provided to the model and fixed to the last vode by overwriting its state $h$. Note that, since both during training and evaluation the state of the first vode would be fixed to the input $x$, we avoid defining it (i.e., we avoid computing $P_{\theta_{0}}(h_{0})$ since it would be constant), and directly forward $x$ to the first layer transformation.
@@
  • JAX [Bradbury et al., 2018] is a functional library, PCX is not. Modules in PCX are PyTrees, using the same philosophy as another popular JAX library, equinox [Kidger and Garcia, 2021], with which PCX modules are fully compatible. However, their state is managed by PCX so that each parameter transformation is automatically tracked. The user can opt in for this behavior by passing arguments as keyword argmunets (such as in the above example). Positional function parameters, instead are ignored by PCX and it is the user’s duty to track their state as done in JAX or equinox.
@@ Appendix B Discriminative experiments
For each model, we conducted experiments with seven different algorithms:
@@
    Standard PC with Cross-Entropy Loss (PC-CE) / Squared Error Loss (PC-SE): already discussed in the background section.
@@
    Unlike standard Predictive Coding with Squared Error Loss (PC-SE), where the output is clamped to the target, we “nudge” the output towards the target in PC with nudging. This is achieved by fixing the representation $h$ of last layer $h_{L}$ to $\mu_{L}+\beta(y-\mu_{L})$, where $\mu_{L}$ is the predicted activation of the last layer after forward initialisation, $y$ is the target, and $\beta\in(0,1)$ is a scalar parameter that controls the strength of nudging. Note that when $\beta=1$, PC with nudging is equivalent to the standard PC.
    During training procedure, as the model output gradually approaches to the target, we employ a strategy of increasing $\beta$. At the end of each epoch, the value of $\beta$ is incremented by a fixed learning rate $\beta_{ir}$. When $\beta$ becomes greater than or equal to 1, we set it to 1. This strategy allows the model to learn and explore in the early stages of training, while gradually transitioning to standard PC in the later stages.
@@
    In this algorithm, we do the opposite of positive nudging: we push the output away from the target. Therefore, we fix the representation $h$ of the last layer to $\mu_{L}-\beta(y-\mu_{L})$. We use the same strategy of dynamically increasing $\beta$. When $\beta$ becomes greater than or equal to 1, we set it to -1.
    In the learning stage, to ensure that the direction of the weight update is consistent with the target (since we fixed $h_{L}$ to the opposite direction), we invert the weight update: $\theta_{l}\leftarrow\theta_{l}-\Delta\theta_{l}$ where $\Delta\theta_{l}$ defined in the Eq. 3.
@@
    Center Nudging [Scellier et al., 2024] is used in equilibrium propagation to improve and stabilize performance compared to both positive and negative nudging, and it is obtained as an average of the gradients produced by the two methods. Here, we approximate this behavior by randomly alternating between epochs in which we train with either negative or positive nudging. In this way, the training model can benefit from both methods without any extra computational cost.
@@
    Incremental PC (iPC), a simple and recently proposed modification where the weight parameters are updated alongside the latent variables at every time step [Salvatori et al., 2024].
@@
    Standard Backpropagation with Cross-Entropy Loss (BP-CE) / Squared Error Loss (BP-SE): the most popular way to do the credit assignment in the neural networks. The model is trained by computing the gradients of the loss function with the weights of the network using the chain rule.
@@
The model hyperparameters are determined using the search space shown in Table 7. The results presented in Table 1 were obtained using 5 seeds with the optimal hyperparameters, which are stored in the YAML files located in the "examples/s4_1_discriminative_mode/" subdirectories of the PCX library.
As for the optimizer and scheduler, we use mini-batch gradient descent (SGD) with momentum as the optimizer for the $h$, and we utilize AdamW with weight decay as the optimizer for the $\theta$. Additionally, we apply a warmup-cosine-annealing scheduler without restart for the learning rates of $\theta$. We also tried SGD with momentum for the weights $\theta$, but we did not perform complete hyperaparemter searches on all combinations of architectures and datasets as its performance was suboptimal compared to AdamW in all the cases tested.
@@ Results.
Activation function also plays a crucial role in improving model accuracy. For models using Cross-Entropy Loss, the “HardTanh” activation function is a better choice. In the case of models using Squared Error Loss without nudging, the “LeakyReLU” activation function tends to perform better. When using Positive Nudging, the optimal activation function varies depending on the model architecture. For Negative Nudging, the “GeLU” activation function is the most suitable choice.
@@
Figure 10: Different training runs for VGG-7 on TinyImageNet using CN. Most of the seeds diverge, resulting in poor average performance.
Training instability and model size. Similarly to what we show in Section 5.1, we noticed that largest architectures present significantly more training instability. iPC is stable and produces optimal results only for the smallest, fully connected, architectures, while PC achieves its peak performance on VGG-5. In particular, we noticed that VGG-7 is able to reach similar highest accuracy (e.g., $\approx 43\%$ on Tiny ImageNet) but, on average, performs notably worse, as most seeds result in a diverged model (Fig. 10).
@@ C.1 Autoencoder
Figure 12: Left: An Autoencoder implemented with backpropagation consists of both an encoder and a decoder. The encoder compresses the input data into the bottleneck dimension, and the decoder restores the original image. Right: An Autoencoder implemented with Predictive Coding. The state of the first PC layer is the bottleneck dimension. The state of the last PC layer is the original input, and the predicted state of the last PC layer is the predicted input. Inference steps update the bottleneck dimension to make it a good compressed representation.
@@ C.2 MCPC
MCPC’s implementation in PCX utilizes a noisy SGD optimizer for the state $h$. Compared to PC than uses an SGD or Adam optimizer, MCPC incorporates an optimizer that merges the addition of noise to the model’s gradients with an SGD optimizer. The variance of the noise added to the gradients needs to be carefully crafted to scale appropriately with the learning rate and the momentum as shown in equations (5 - 7).
@@ C.3 Associative memories
Here, the benchmark results are obtained with Tiny ImageNet, corrupted with either Gaussian noise with 0.2 standard deviation, or a mask on the bottom half of the images (examples shown in Fig. 4). We vary the model size and number of training examples to memorize, to study the capacity of the models. Specifically, we use a generative PCN with architecture $[512,d,d,12288]$ where $d=[512,1024,2048]$ ($12288$ being the flattened Tiny ImageNet images) and varied $n=[50,100,250]$. We performed a hyperparameter search for each $d$ and $n$ on the parameter learning rate $lr_{\theta}\in\{1\times 10^{-4}+k\cdot 5\times 10^{-5}\mid k\in\mathbb{Z},0\leq n\leq 18\}$, the state learning rate $\gamma\in\{0.1+k\cdot 0.05\mid k\in\mathbb{Z},0\leq n\leq 18\}$, training inference steps $T_{\textnormal{train}}\in[20,50,100]$ and recall inference steps $T_{\textnormal{recall}}\in[50000,100000]$. We fix the activation function of the model to Tanh, and the number of training epochs to $500$ and a batch size of $50$. The results in Table 3 are obtained with $5$ seeds with the searched optimal hyperparameters, which are stored in the hps.json file under the examples/s4_2_generative_mode/associative_memory folder in the PCX library.
@@ Appendix D Energy and Stability
This section describes the experimental setup of Section 5.1, provides replications on other datasets and ablations.
The code can be found in: examples/s5_analysis_and_metrics/energy_and_stability.
@@ Results.
Fig. 5(b) in the main paper shows the average energy across the last batch at the end of training for the best performing model on the grid. Fig. 5(c) compares SGD with momentum $0.9$ and AdamW. It is obtained for activation function “HardTanh” and a width of 1024. We replicate this figure for the other combinations of activation functions and widths below in Fig 14. We observe that across all conditions, small to medium state learning rates are generally preferred by SGD, while AdamW has a stronger preference to smaller state learning rates. Given the uneven distribution of energies across layers, Adam, in particular, may not scale to deeper architectures. We further, observe a larger variance in performance for Adam, especially for wider layers, which we discuss in paragraph “Training Instability“ in Sec. 5.1 and below. Fig. 5(d) is based on all models trained with AdamW. Many models with high state learning rates diverge, we only plot models achieving accuracy $>0.5$.
@@
Figure 15: Energy propagation on the Two Moons dataset. 15(a) shows the imbalance between layers across $T$ steps. 15(b) shows the model performance across state learning rates and 15(c) the energy distribution across state learning rates.
Figure 16: Energy propagation on the Two Circles dataset. 16(a) shows the imbalance between layers across $T$ steps. 16(b) shows the model performance across state learning rates and 16(c) the energy distribution across state learning rates.
Below we present the results of experiments on the Two Moons and Two Circles datasets. Fig. 15(b), 15(a), and 15(c) replicate figures Fig. 5(b),  5(c), and 5(d) for Two Moons, and Fig. 16(b), 16(a), and 16(c) for Two Circles. Results are very similar to FashionMNIST: The energy is concentrated in the last layer, even after $T$ inference steps. However, in the example for Two Circles, we actually observe a training effect for earlier layers: While the energy increases first due to error propagation (still orders of magnitude below later layers), the energy is reduced afterwards. Energy ratios are consistenly indicating poor energy propagation for state learning rates $\gamma$, that perform well. As predicted the variance in results is significantly larger for Two Circles, especially for small state learning rates.
@@
We replicate Fig. 6 (FashionMNIST) here for the Two Moons dataset, see Fig. 17. We observe effects for Two Moons that are analog to FashionMNIST as presented above: The stability of optimization strongly depends on the width of the hidden layers for Adam. This effect is not observed for SGD on either dataset. This further supports the our conclusion in Sec. 5.1: While Adam is the better optimizer, this interaction effect (width $\times$ $\gamma$) can hinder the scaling of PCNs with Adam. Optimization methods for PCNs require further attention from the research community.
@@
We further provide an ablation on FashionMNIST. In the experiments above, the hidden layer width is altered, introducing changes in the absolute size of the hidden layers (i.e. number of neurons), but also changing the relative size of the hidden layers in the network, as input and output layers remain the same size across all experiments. Hence, we provide another experiment on FashionMNIST, where we increase the image size and augment the label vector with 00s, such that the width of all layers is equal. All other experimental variables remain as described above. The results are shown in Fig. 18 and follow the trend observed in Fig. 6 and 17: We find that there exists an interaction between the optimization and the width of the network as described above. Hence, accounting for relative changes in layer width does not sufficiently explain the problem and we conclude that the absolute size of the layers plays a role in the stability of optimization with AdamW.
D.3 Integrating Skip Connections into VGG19 to Enhance PC Networks for CIFAR10 Classification

### v2

@@ 6 Library, Resources and Implementations Details
The efficiency of PCX could be further increased by fully parallelizing all the operations. In fact, in its current state, JIT is unable to parallelize the execution of the layers; a problem that can be addressed with the JAX primitive vmap, but only in the unpractical case where all the layers have the same dimension. To test how different hyperparameters of the model influence the training speed, we have taken a feedforward model, and trained it multiple times, each time increasing a specific hyperparameter by a multiplicative factor. The results, reported in Fig. 8, show that the two parameters that increase the training time are the number of layers $L$ and the number of steps $T$. Ideally, only $T$ should affect the training time as inference is an inherently sequential process that cannot be parallelized, but this is not the case, as the time scales linearly with the amount of layers. Details are reported in Appendix G.
@@ 7 Discussion
The main contribution of this work is the introduction and open-source release of PCX, a library that can be used to perform deep learning tasks using PCNs. Its efficiency relies on JAX’s Just-In-Time compilation and carefully structured primitives built to take advantage of it. A second advantage of our library is its intuitive setup, tailored to users already familiar with other deep learning frameworks such as PyTorch. This, together with the large number of tutorials we release, will make it easy for new users to train networks using PC. We have then used PCX to perform an extensive comparative study among different models and training algorithms present in the literature, obtained by testing a large number of parameter combinations and activation functions.
In terms of results, we have shown that predictive coding networks perform comparably to standard deep learning ones trained with BP, conditioned on the fact that small/medium size architectures are used, such as VGG 7. When this condition is relaxed, the performance of predictive coding fails to match that of BP, able to scale along with model size. In the supplementary material, we add rigorous studies that provide more details about how the energy flows inside PCNs over time, and their training stability, as well as show how PCNs classify out-of-distribution data, and possible solutions for training extremely deep networks via the use of skip connections.
@@ Acknowledgments
We thank the reviewers for their valuable feedback and insightful discussions, which have significantly enhanced this manuscript. Amine M’Charrak gratefully acknowledges support from the Evangelisches Studienwerk e.V. Villigst through a doctoral fellowship. Tommaso Salvatori gratefully acknowledges funding from VERSES AI. Rafal Bogacz gratefully acknowledged support by Medical Research Council grant MC_UU_00003/1. Cornelius Emde is supported by the EPSRC Centre for Doctoral Training in Health Data Science (EP/S02428X/1). Thomas Lukasiewicz and Cornelius Emde are supported by the AXA Research Fund.
@@ Appendix A PCX – A Brief Introduction
In Section 3, we defined PCNs as models with parameters $\theta=\{\theta_{0},\ldots,\theta_{L}\}$ and state $h=\{h_{0},\ldots,h_{L}\}$. In PCX, we divide a model in two main components: layers (i.e., the traditional deep-learning transformations such as ’Linear’ or ’Conv2D’) and vodes (i.e., vectorized nodes that store the array of neurons representing state $h_{l}$). A PCN is defined as follows: {minted}python import jax.nn as jnn import pcx.predictive_coding as pxc import pcx.nn as pxnn
@@
return u In the __call__ method, we forward the input $x$ through the network. Note that every time we call a vode, we are effectively storing in it the activation $u_{l}$ (so that we can later compute the energy $\epsilon_{l}^{2}$ associated to the vode) and return its state $h_{l}$ (i.e., x = vode(u) corresponds to vode.set("u", u); x = vode.get("h")). During training, the label $y$ is provided to the model and fixed to the last vode by overwriting its state $h^{(L)}$. Note that, since both during training and evaluation the state of the first vode would be fixed to the input $x$, we avoid defining it (i.e., we avoid computing $P_{\theta_{0}}(h_{0})$ since it would be constant), and directly forward $x$ to the first layer transformation.
@@
  • JAX (Bradbury et al., 2018) is a functional library, PCX is not. Modules in PCX are PyTrees, using the same philosophy as another popular JAX library, equinox (Kidger & Garcia, 2021), with which PCX modules are fully compatible. However, their state is managed by PCX so that each parameter transformation is automatically tracked. The user can opt in for this behavior by passing arguments as keyword arguments (such as in the above example). Positional function parameters, instead are ignored by PCX and it is the user’s duty to track their state as done in JAX or equinox.
@@ Appendix B Discriminative experiments
For each model, we conducted experiments with the following different algorithms:
@@
    Standard PC with Cross-Entropy Loss (PC-CE) / Mean Squared Error Loss (PC-SE): already discussed in the background section.
@@
    Unlike standard Predictive Coding with Mean Squared Error Loss (PC-SE), where the output is clamped to the target, we “nudge” the output towards the target in PC with nudging. This is achieved by fixing the representation $h$ of last layer $h_{L}$ to $\mu_{L}+\beta(y-\mu_{L})$, where $\mu_{L}$ is the predicted activation of the last layer after forward initialisation, $y$ is the target, and $\beta\in(0,1)$ is a scalar parameter that controls the strength of nudging. Note that when $\beta=1$, PC with nudging is equivalent to the standard PC.
    During training procedure, as the model output gradually approaches to the target, we employ a strategy of increasing $\beta$. At the end of each epoch, the value of $\beta$ is incremented by a fixed rate $\beta_{ir}$. When $\beta$ becomes greater than or equal to 1, we set it to 1. This strategy allows the model more stable to learn and explore in the early stages of training, while gradually transitioning to the standard PC in the later stages.
@@
    In this algorithm, we do the opposite of positive nudging: we push the output away from the target. Therefore, we fix the representation $h$ of the last layer to $\mu_{L}-\beta(y-\mu_{L})$. We use the same strategy of dynamically increasing $\beta$. When $\beta$ becomes greater than or equal to -1, we set it to 1.
    In the learning stage, to ensure that the direction of the weight update is consistent with the target (since we fixed $h_{L}$ to the opposite direction), we invert the weight update: $\theta_{l}\leftarrow\theta_{l}-\Delta\theta_{l}$ where $\Delta\theta_{l}$ defined in the Eq. (3).
@@
    Center Nudging (Scellier et al., 2024) is used in equilibrium propagation to improve and stabilize performance compared to both positive and negative nudging, and it is obtained as an average of the gradients produced by the two methods. Here, we approximate this behavior by randomly alternating between epochs in which we train with either negative or positive nudging. In this way, the training model can benefit from both methods without any extra computational cost.
@@
    Incremental PC (iPC), a simple and recently proposed modification where the weight parameters are updated alongside the latent variables at every time step (Salvatori et al., 2024).
@@
    Standard Backpropagation with Cross-Entropy Loss (BP-CE) / Mean Squared Error Loss (BP-SE): the most popular way to do the credit assignment in the neural networks. The model is trained by computing the gradients of the loss function with the weights of the network using the chain rule.
@@
The model hyperparameters are determined using the search space shown in Table 7. The results presented in Table 1 were obtained using 5 seeds with the optimal hyperparameters.
As for the optimizer and scheduler, we use mini-batch gradient descent (SGD) with momentum as the optimizer for the $h$, and we utilize AdamW Loshchilov & Hutter (2017) with weight decay as the optimizer for the $\theta$. Additionally, we apply a warmup-cosine-annealing scheduler without restart for the learning rates of $\theta$.
@@ Results.
Activation function also plays a crucial role in improving model accuracy. For models using Cross-Entropy Loss, the “HardTanh” activation function is a better choice. In the case of models using Mean Squared Error Loss without nudging, the “LeakyReLU” activation function tends to perform better. When using Positive Nudging, the optimal activation function varies depending on the model architecture. For Negative Nudging, the “GeLU” activation function is the most suitable choice.
@@ C.1 Autoencoder
Figure 11: Left. An Autoencoder implemented with backpropagation consists of both an encoder and a decoder. The encoder compresses the input data into the bottleneck dimension, and the decoder restores the original image. Right. An Autoencoder implemented with Predictive Coding. The state of the first PC layer is the bottleneck dimension. The state of the last PC layer is the original input, and the predicted state of the last PC layer is the predicted input. Inference steps update the bottleneck dimension to make it a good compressed representation.
@@ C.2 MCPC
MCPC’s implementation in PCX utilizes a noisy SGD optimizer for the state $h$. Compared to PC than uses an SGD or Adam optimizer, MCPC incorporates an optimizer that merges the addition of noise to the model’s gradients with an SGD optimizer. The variance of the noise added to the gradients needs to be carefully crafted to scale appropriately with the learning rate and the momentum as shown in equations (4 - 6).
@@ C.3 Associative memories
Here, the benchmark results are obtained with Tiny ImageNet, corrupted with either Gaussian noise with 0.2 standard deviation, or a mask on the bottom half of the images (examples shown in Fig. 5). We vary the model size and number of training examples to memorize, to study the capacity of the models. Specifically, we use a generative PCN with architecture $[512,d,d,12288]$ where $d=[512,1024,2048]$ ($12288$ being the flattened Tiny ImageNet images) and varied $n=[50,100,250]$. We performed a hyperparameter search for each $d$ and $n$ on the parameter learning rate $lr_{\theta}\in\{1\times 10^{-4}+k\cdot 5\times 10^{-5}\mid k\in\mathbb{Z},0\leq n\leq 18\}$, the state learning rate $\gamma\in\{0.1+k\cdot 0.05\mid k\in\mathbb{Z},0\leq n\leq 18\}$, training inference steps $T_{\textnormal{train}}\in[20,50,100]$ and recall inference steps $T_{\textnormal{recall}}\in[50000,100000]$. We fix the activation function of the model to Tanh, and the number of training epochs to $500$ and a batch size of $50$. The results in Table 3 are obtained with $5$ seeds with the searched optimal hyperparameters.
@@ Appendix D Energy and Stability
This section describes the experimental setup of Section LABEL:sec:energy-stability, provides replications on other datasets and ablations.
@@ Results.
Fig. 6(left) in the main paper shows the average energy across the last batch at the end of training for the best performing model on the grid. Fig. 6(center-left) compares SGD with momentum $0.9$ and AdamW. It is obtained for activation function “HardTanh” and a width of 1024. We replicate this figure for the other combinations of activation functions and widths below in Fig 13. We observe that across all conditions, small to medium state learning rates are generally preferred by SGD, while AdamW has a stronger preference to smaller state learning rates. Given the uneven distribution of energies across layers, AdamW, in particular, may not scale to deeper architectures. We further, observe a larger variance in performance for AdamW, especially for wider layers, which we discuss in paragraph “Training Instability“ in Sec. LABEL:sec:energy-stability and below. Fig. 6(right) is based on all models trained with AdamW. Many models with high state learning rates diverge, we only plot models achieving accuracy $>0.5$.
@@
Figure 14: Energy propagation on the Two Moons dataset. 14(a) shows the imbalance between layers across $T$ steps. 14(b) shows the model performance across state learning rates and 14(c) the energy distribution across state learning rates.
Figure 15: Energy propagation on the Two Circles dataset. 15(a) shows the imbalance between layers across $T$ steps. 15(b) shows the model performance across state learning rates and 15(c) the energy distribution across state learning rates.
Below we present the results of experiments on the Two Moons and Two Circles datasets. Fig. 14(b), 14(a), and 14(c) replicateFig. 6 for Two Moons, and Fig. 15(b), 15(a), and 15(c) for Two Circles. Results are very similar to FashionMNIST: The energy is concentrated in the last layer, even after $T$ inference steps. However, in the example for Two Circles, we actually observe a training effect for earlier layers: While the energy increases first due to error propagation (still orders of magnitude below later layers), the energy is reduced afterwards. Energy ratios are consistenly indicating poor energy propagation for state learning rates $\gamma$, that perform well. As predicted the variance in results is significantly larger for Two Circles, especially for small state learning rates.
@@
We replicate Fig. 7 (FashionMNIST) here for the Two Moons dataset, see Fig. 16. We observe effects for Two Moons that are analog to FashionMNIST as presented above: The stability of optimization strongly depends on the width of the hidden layers for Adam. This effect is not observed for SGD on either dataset. This further supports the our conclusion in Sec. LABEL:sec:energy-stability: While Adam is the better optimizer, this interaction effect (width $\times$ $\gamma$) can hinder the scaling of PCNs with Adam. Optimization methods for PCNs require further attention from the research community.
@@
We further provide an ablation on FashionMNIST. In the experiments above, the hidden layer width is altered, introducing changes in the absolute size of the hidden layers (i.e. number of neurons), but also changing the relative size of the hidden layers in the network, as input and output layers remain the same size across all experiments. Hence, we provide another experiment on FashionMNIST, where we increase the image size and augment the label vector with 00s, such that the width of all layers is equal. All other experimental variables remain as described above. The results are shown in Fig. 17 and follow the trend observed in Fig. 7 and 16: We find that there exists an interaction between the optimization and the width of the network as described above. Hence, accounting for relative changes in layer width does not sufficiently explain the problem and we conclude that the absolute size of the layers plays a role in the stability of optimization with AdamW.
ResNets
Here we discuss the findings on the energy propagation in light of the ResNets18 experiments. In this section, we have shown that lower learning rate for the nodes harm energy propagation, and that the AdamW optimizer displays poor performance for larger hidden dimensions. To this end, we have trained ResNets18 using SGD and large learning rates for the nodes, and compared the performance against those in the main body of the paper. The performance are, however, not comparable to the ones reported in Table.1, as ResNets trained with SGD on the CIFAR10 dataset reach accuracies of $39.9\%$ and $43.2\%$ when using PC and iPC, respectively. To better understand the incidence of different hyperparameters on the final test accuracy of the models, in Fig. 18 we show their importance plots. Such quantities are computed by fitting a random forest regressor with hyperparameters as datapoints, accuracies as labels, and extracting the feature importance.
Figure 18: Importance plots that show the importance of each hyperparameter in the final test accuracy of the model, computed by fitting a random forest regressor with hyperparameters as datapoints, accuracies as labels, and extracting the feature importance.
Appendix E Skip Connections into VGG19